# Rubrics as Rewards: Reinforcement Learning Beyond Verifiable Domains

### v1

@@ 4.5 Evaluation Setup
Medical Reasoning
We evaluate models trained on curated medical reasoning data using HealthBench-1k Arora et al. (2025), a 1,000-example subset of HealthBench. [^3] We introduce HealthBench-1k as a cost-effective evaluation split for medical reasoning tasks and will release the test set identifiers to support reproducibility.. We generate model responses with greedy decoding (temperature=0) and report the overall_score.
Science Reasoning
We evaluate models trained with curated science set on the GPQA-Diamond subset Rein et al. (2024). Each evaluation is independently repeated 4 times to account for variance, where each run samples one response per prompt using greedy decoding (temperature=0). Answer choices are permuted per example, and model outputs are parsed for boxed answer formats (e.g.,
boxed{A}). If extraction fails, we fall back to a GPT-4o verifier that checks whether the response contains the correct option letter or text (see F). Final accuracy is reported as the mean of the 4 independent runs.
@@ 5 Results
Rubrics as Rewards Combine Interpretability with Strong Performance.
Table 1 reports results on the Medicine (HealthBench-1k) and Science (GPQA_Diamond) benchmarks. Our method, RaR-Implicit, consistently outperforms baselines such as Simple-Likert, with the best variant achieving upto 28% relative improvement on HealthBench-1k and 13% on GPQA. It also surpasses both base and instruction-tuned policy models, demonstrating the effectiveness of rubric-guided training for nuanced response evaluation.
Furthermore, RaR-Implicit matches or exceeds the performance of the Reference-Likert baseline. In our setup, synthetic rubrics are generated using reference answers as proxies for expert supervision, meaning their quality is inherently constrained by the completeness and clarity of those references. Despite this, transforming open-ended answers into structured rubric criteria produces reward signals that are both effective and aligned with evaluation objectives.
We also find that implicit aggregation outperforms explicit, weighted summation which indicates that allowing an LLM-based judge to infer how to balance rubric criteria per prompt is more effective than relying on fixed, hand-crafted weights. While explicit aggregation offers more transparency and control, implicit methods better adapt to real-world complexity.
Rubric-based supervision further enables transparent reward specification by breaking response quality into interpretable criteria. Unlike holistic reference comparisons, rubrics reduce stylistic bias and surface-level overfitting. Though reference-based Likert scoring may perform slightly better in some cases, rubrics offer a more scalable and controllable approach—particularly where references are scarce or model interpretability is essential.
Figure 2: LLM Judge Alignment Across Model Scales. Rubric-guided evaluation (orange) consistently improves judge accuracy across model sizes compared to pure Likert-based scoring (blue). Rubrics without reference answers (green) perform better than Likert baseline, though still lag behind full rubric guidance. Smaller scale models benefit the most from rubric structure, narrowing the performance gap with larger judges.
@@
To evaluate the effectiveness of rubric-based evaluations in reflecting human preferences, we generated datasets consisting preferred (chosen) and perturbed (rejected) responses (see Appendix I for details on dataset generation). We use LLM-Judges to assign ratings from 1 to 10 for these responses both with (RaR-IMPLICIT ) and without the guidance of rubrics (SIMPLE-LIKERT). Figure 2 shows how often judges correctly rated the preferred responses higher than the rejected ones across various model scales (from 3b to 72b to o3-mini). The results consistently show that rubric-guided evaluations provide clearer and more accurate signals, as evidenced by the higher accuracy with which preferred responses receive higher ratings when rubrics are involved. This indicates that rubrics improves the alignment with human preferences by offering detailed, context-specific criteria that help judges better discern subtle differences in quality compared to evaluations using only simple Likert-based ratings. Further study on LLM-Judge scaling impacts on policy training are discussed in Appendix C.
@@ 6 Ablations
This ablation study examines how the structure and weighting of synthetic rubrics—generated using o3-mini—affect downstream performance on HealthBench. As shown in Table 2, predefined generic rubrics substantially underperform compared to prompt-specific ones, underscoring the importance of contextualization. Rubrics that include a broader range of criteria—both positive and negative—consistently outperform those limited to essential checks, suggesting that richer evaluation signals lead to better learning. Interestingly, we observe minimal performance differences when including rubric weights or pitfall criteria during training. One possible explanation is that synthetically generating effective pitfall criteria is inherently difficult, as it requires anticipating the most common or critical failure modes of the model, a task that often demands human intuition and domain expertise. As a result, these synthetic negative criteria may lack the specificity or relevance needed to meaningfully penalize undesirable responses.
Ablation Setting Overall Score
Predefined Generic Rubrics 0.2898 $\pm$ 0.00526
Essential-Only Rubrics 0.3562 $\pm$ 0.00976
No Weights 0.3756 $\pm$ 0.00479
No-Pitfall Criteria 0.3837 $\pm$ 0.00420
All Rubrics 0.3723 $\pm$ 0.00005
Table 2: Ablation results on HealthBench-1k (trained on HealthBench-3.5k subset with Qwen2.5-7B base policy). Rubrics generated using o3-mini with access to reference answers. Variations test the impact of rubric structure and weighting on implicit reward aggregation.
Training Method Overall Score
Expert-Answer-SFT 0.204
Simple-Likert 0.239
Reference-Likert 0.317
RaR-Implicit-Synthetic-NoRef 0.320
RaR-Implicit-Synthetic 0.359
RaR-Implicit-Human 0.348
Impact of Rubric Generation Strategies in Real-World Domains
How does the method of rubric generation affect downstream training in challenging, real-world settings? To investigate this, we evaluate on the HealthBench dataset using a curated subset of 3.5k human-annotated examples, split into train and test sets. Table 3 summarizes the results.
We observe that all rubric-based methods outperform rubric-free baselines, including Reference-Likert. Notably, even the weakest RaR variant significantly surpasses Reference-Likert, underscoring the advantage of structured supervision in subjective, open-ended domains like healthcare. We attribute this to the finer granularity and clarity rubrics provide in assigning rewards-especially when correctness is not binary and answers may vary in tone, completeness, or safety relevance.
@@
Rubric Generation Model Overall Score
O3-mini (Rubrics with reference) 0.359
GPT-4o 0.342
GPT-4o-mini 0.327
O3-mini 0.324
Qwen-72B-Instruct 0.327
Qwen-32B-Instruct 0.311
Qwen-7B-Instruct 0.319
@@ 7 Related Work
Reinforcement learning with verifiable rewards (RLVR) is quickly extending its reach well beyond math and code. General-Reasoner trains on a 200 k mixed corpus-physics, finance, policy, and more-and reports a ten-point jump on MMLU-Pro after GRPO fine-tuning Ma et al. (2025). A follow-up study extends RLVR to medicine, chemistry, psychology, and economics, showing that a single cross-domain reward model can supervise all four areas without task-specific tweaks Su et al. (2025a). Focusing on healthcare, Med-RLVR applies the same recipe to multiple-choice clinical QA, gaining eight accuracy points over supervised baselines while eliciting chain-of-thought reasoning from a 3B-parameter base model Zhang et al. (2025). These studies collectively show promising progress in taking RL beyond math and code, yet sparse reward signals, unreliable verifiers, and limited benchmark coverage remain open challenges.
Rubric-based Evaluation
Concise, task-specific rubrics are now standard for evaluating frontier LLMs on open-ended benchmarks (Arora et al., 2025; Ruan et al., 2025; Hashemi et al., 2024; Pathak et al., 2025). Pathak et al. show that rubric-prompted GPT-4 graders are more accurate and consistent than a single question-agnostic checklist (Pathak et al., 2025), yet their rubrics appear only at evaluation time, not in training. HealthBench extends the idea to medicine, pairing 48 k clinician-written criteria with GPT-4 judges to score factuality, safety, and empathy (Arora et al., 2025). Moving from evaluation to tuning, Configurable Preference Tuning (CPT) synthesizes rubric-conditioned preference pairs for DPO fine-tuning (Gallego, 2025). None of these approaches, however, integrates rubrics with verifiable rewards. Rubrics are used only to grade held-out outputs or to generate preference pairs.
@@
RLHF tunes LLMs with thousands of human preference comparisons, but these subjective labels introduce noise and invite reward hacking (Ouyang et al., 2022). Reinforcement learning with verifiable rewards (RLVR) alleviates these issues by relying on programmatic checks—ranging from exact string matches on GSM8K and MATH to mixed-domain verifiers in General-Reasoner and Cross-Domain RLVR Ma et al. (2025); Su et al. (2025a). Although the resulting signals are low-variance and fully automatic, they can be sparse. Process supervision provides more detailed feedback by rewarding intermediate reasoning steps. MCTS-generated labels and generative reward models such as ThinkPRM improve performance, but at a significant annotation cost (Li et al., 2025; Khalifa et al., 2025). Our rubric based RLVR closes this gap by turning rubric criteria into structured verifiers and using their scalar scores as dense rewards in on-policy RLVR training.
8 Future Work
Interpretability and Reward Robustness
RaR offers a transparent alternative to black-box reward models by explicitly optimizing against human-interpretable rubric criteria. Unlike neural reward models or preference-based methods such as DPO, RaR enables auditability and objective attribution of reward signals. This transparency may also increase robustness to reward hacking, as discrete, interpretable criteria are harder for models to exploit than opaque learned scores.
Future work can explore formal evaluations of RaR’s resilience to reward hacking, including adversarial training scenarios where models attempt to game individual rubric components.
Extending to Real-World Domains and Use Cases
Rubric-guided post-training has been underexplored in open-ended or agentic tasks. A promising direction is expanding RaR to domains with subjective or multi-step goals such as tool use, reasoning under uncertainty, or real-world decision support-where rubrics can provide structured, task-specific guidance.
Rubrics as Curriculum
Rubrics naturally encode hierarchical task structure—from essential facts to optional stylistic preferences—which can support implicit curriculum learning. Early in training, models may only satisfy simpler criteria; as learning progresses, more complex targets become attainable. Future work can investigate dynamic weighting or staged introduction of rubric items to further leverage this curriculum effect.
@@ 9 Conclusion
Limitations
Our study is limited to two domains (medical and science), and further validation is needed to assess generalizability to other tasks, such as open-ended dialogue. We explore only two reward aggregation strategies—implicit and explicit—and do not investigate alternative weighting schemes. Additionally, we do not conduct controlled analyses of reward hacking risks. Finally, our judges are off-the-shelf LLMs; future work could explore dedicated evaluators with stronger reasoning capabilities.
@@ Acknowledgments
Appendix A Training dataset breakdown
In this section, we show the topic distribution breakdown of the medicine and science training dataset and the rubric type distribution breakdown for the corresponding dataset.
A.1 Medicine training dataset - Overall Data Distribution
@@
Important 52 748 34.1
Essential 47 584 30.7
Optional 34 261 22.1
Pitfall 20 215 13.1
@@
Total examples 20 166 100.0
Medical Diagnosis 10 147 50.3
Medical Treatment 3 235 16.0
Medical Knowledge 2 557 12.7
Medical Diag. and Mngmnt 2 033 10.1
@@
A.2 STEM Dataset - Overall Data Distribution

### v2

@@ 4.5 Evaluation Setup
Rubric-Based Evaluation
We evaluate models trained with RaR-Medicine on HealthBench [2], a benchmark of 5,000 clinical conversations designed to assess model safety and helpfulness in realistic medical scenarios. Performance is measured using detailed, physician-authored rubrics. We generate responses with greedy decoding (temperature = 0) and report both overall scores and per-axis scores following the original setup. For ablation studies, we sample a subset of 1,000 prompts (hereafter referred as HealthBench-1k) and use the rest for training.
Multiple-Choice Evaluation
Each model is evaluated across 10 independent runs, using greedy decoding (temperature=0) to sample one response per prompt. Answer choices are permuted per example to reduce positional bias, and outputs are parsed for boxed answer formats (e.g., boxed{A}). If extraction fails, we fall back to a GPT-4o verifier that checks whether the response contains the correct option letter or text (see Appendix A.4). Final accuracy is reported as the mean over 10 runs, and we include 95% confidence intervals to account for run-to-run variance.
Figure 2: Performance of baselines and RaR (Rubrics as Rewards) variants for the medicine and science domains. HealthBench (left): shows per-axis scores across five core axes, with a thin dashed gray line indicating the overall score (all values shown as percentages). GPQA-Diamond (right): mean accuracy over 10 runs; error bars represent 95% confidence intervals. All policies are evaluated using gpt-4o-mini as the LLM-as-Judge. Across both domains, RaR-Implicit consistently outperforms Direct-Likert and demonstrates a competitive advantage over Reference-Likert.
LLM-Judge Alignment Evaluation
To measure how well LLM judges align with human preferences, we build a paired evaluation set from roughly 3,000 HealthBench prompts. For each prompt, we take the practitioner-approved answer as the preferred response and create a perturbed alternative via controlled edits (see Appendix A.9 for method used for perturbation and prompt selection). The metric is pairwise preference accuracy i.e. the fraction of pairs where the preferred response scores higher reported across judge models of varying sizes.
@@ 5 Results
Rubrics as Rewards shows strong gains across evaluation settings.
Table 2 reports results on HealthBench (rubric-based, free-form) and GPQA-Diamond (multiple-choice). RaR-Implicit consistently outperforms Direct-Likert, with relative gains up to 31% on HealthBench and 7% on GPQA. Both rubric-guided variants achieve higher scores than the base and instruction-tuned policies. Gains on GPQA-Diamond show that rubric-induced skills generalize beyond rubric-based evaluation. The RaR-Predefined variant, which applies a fixed list of generic rubrics to every prompt (no instance-specific synthesis), underperforms because generic criteria miss prompt-specific requirements and common failure modes, producing misaligned reward signals. Hence, effective training requires instance-specific rubric synthesis as they better capture task context and typical failure modes.
Beyond these gains, RaR-Implicit also shows small but consistent gains over Reference-Likert. In our setup, rubrics are generated with stronger LLMs using reference answers as proxies for expert supervision, so rubric quality is impacted by reference quality. Even so, converting open-ended answers into explicit criteria yields effective, well-aligned reward signals.
Between the two rubric-guided methods, RaR-Implicit attains the strongest results overall; fixed weighted sums in RaR-Explicit offer more control but can be brittle. Explicit weighting can be difficult to tune but offers greater interpretability; we view the choice as application-dependent and leave it to practitioners. Future work could explore learned or dynamic weighting strategies that maintain interpretability while improving adaptability.
Figure 3: Alignment Study of LLM Judges across Model Scales. Rubrics as Rewards (orange) consistently improves alignment with human preferences across LLM judge sizes compared to direct Likert-based scoring (blue). Judge Alignment using synthetic rubrics without expert grounding (green) outperform the direct Likert baseline, but still fall short of expert-grounded rubrics (orange). The rubric structure especially benefits smaller judge models, helping them close the gap with larger models when guided by checklist-style criteria.
@@
We evaluate alignment with humans by having LLM judges of varying sizes score chosen vs. rejected HealthBench-1k responses on a 1–10 scale under two settings: (i) rubric-guided (RaR-IMPLICIT), where the instance-specific rubric is provided, and (ii) rubric-free (DIRECT-LIKERT), where only the prompt and answers are shown. Figure 3 reports pairwise preference accuracy (the fraction of pairs where the preferred response receives the higher score). Rubric guidance improves accuracy for every judge size, with the largest gains for smaller judges, narrowing the gap to larger models. This indicates that explicit, context-specific criteria help judges distinguish subtle quality differences better than direct Likert scoring. Further analysis of judge-scale effects on GRPO training is detailed in Appendix A.8.
@@
Training Method Overall Score
Expert-Answer-SFT 20.4%
Simple-Likert 23.9%
Reference-Likert 31.7%
RaR-Implicit-Synthetic-NoRef 32.0%
RaR-Implicit-Synthetic 35.9%
RaR-Implicit-Human 34.8%
Ablation Setting Overall Score
Essential-Only Rubrics 34.9%
No Categorical Labels 38.8%
No Pitfall Criteria 37.2%
All Rubrics 37.2%
Table 2: Ablation results for elements of rubric design on HealthBench-1k (trained on HealthBench-3.5k subset with Qwen2.5-7B base policy). Rubrics generated using o3-mini with access to reference answers.
@@ 6 Ablations
Rubric Generation Model Overall Score
O3-mini (Rubrics with reference) 35.9%
GPT-4o 34.2%
GPT-4o-mini 32.7%
O3-mini 32.4%
Qwen-72B-Instruct 32.7%
Qwen-32B-Instruct 31.1%
Qwen-7B-Instruct 31.9%
Impact of Rubric Generation Strategies in Real-World Domains
How does the method of rubric generation affect downstream training in challenging, real-world settings? To investigate this, we hold out HealthBench-1k for evaluation and use 3.5k prompts from the remaining HealthBench pool to generate rubrics for training as it has access to human generated rubrics. Results are summarized in Table 1.
In-domain testing on HealthBench-1k amplifies RaR’s gains: every instance-specific rubric-based method outperforms rubric-free baselines. Notably, even the weakest RaR variant significantly surpasses Reference-Likert, underscoring the advantage of structured supervision in subjective, open-ended domains like healthcare. We attribute this to the finer granularity and clarity rubrics provide in assigning rewards-especially when correctness is not binary and answers may vary in tone, completeness, or safety relevance.
@@
This ablation study examines how the structure and weighting of synthetic rubrics affect downstream performance on HealthBench-1k. As shown in Table 2, rubrics that include a broader range of criteria outperform those limited to essential checks, suggesting that richer evaluation signals lead to better learning. Interestingly, we observe minimal performance differences when including rubric weights or pitfall criteria during training. One possible explanation is that synthetically generating effective pitfall criteria is inherently difficult, as it requires anticipating the most common or critical failure modes of the model, a task that often demands human intuition and domain expertise. As a result, these synthetic negative criteria may lack the specificity or relevance needed to meaningfully penalize undesirable responses.
@@ 7 Related Work
Reinforcement learning with verifiable rewards (RLVR) is expanding beyond math and code. General-Reasoner trains on a 200k mixed corpus spanning physics, finance, and policy, and reports a ten-point gain on MMLU-Pro after GRPO fine-tuning [23]. A follow-up extends RLVR to medicine, chemistry, psychology, and economics, showing that a single cross-domain reward model can supervise all four without task-specific tweaks [31]. In healthcare, Med-RLVR applies similar methods to multiple-choice clinical QA, improving accuracy over supervised baselines while eliciting chain-of-thought from a 3B base model [38]. These results indicate steady progress, yet sparse signals, verifier reliability, and limited benchmark coverage remain open challenges.
Rubrics for evaluation and training
Task-specific rubrics are increasingly used to evaluate LLMs in difficult-to-verify domains [2, 27, 13, 26]. Pathak et al. show that rubric-prompted LLM graders are more accurate and consistent than a question-agnostic checklist [26]. HealthBench scales this idea in medicine, pairing 48k clinician-written criteria with GPT-4 judges to score various axes [2]. Beyond evaluation, rubrics are used to condition preference pairs for DPO (CPT; [8]) and to guide checklist-based preference tuning in safety, instruction-following, and creative-writing settings [33, 7, 18]. These lines of work primarily use rubrics to grade outputs or to condition preference data, often in non-reasoning domains such as safety, instruction following, or creative writing. In contrast, we use rubric criteria directly as reward signals for on-policy RL in expert-reasoning and real-world domains.
@@
RLHF trains policies with large numbers of human comparisons, which introduces subjectivity and can lead to reward hacking [25]. RLVR reduces these issues by using programmatic checks, from exact matches on GSM8K and MATH to mixed-domain verifiers in General-Reasoner and Cross-Domain RLVR [23, 31], although signals can be sparse. Process supervision [21] provides denser guidance via step-level labels, and MCTS-generated annotations or generative reward models such as ThinkPRM improve performance, but with high annotation cost [20, 16]. Rubric-based RL finds a middle ground by turning multiple rubric criteria into structured verifiers and using their scalar scores as denser rewards.
@@ 8 Conclusion
9 Limitations and Future Work
Our work focuses on medicine and science to enable controlled experiments. This choice allows us to run controlled experiments, but broader validation across dialogue, tool use, or other agentic tasks remains an important direction. We evaluate only two reward aggregation strategies, implicit and explicit, since they capture complementary extremes of flexibility and control; future work could explore more advanced ways of combining rubric criteria, such as learning continuous weights for each criterion or dynamically adjusting weights over the course of training to mimic a curriculum (e.g., prioritizing essential correctness early, then gradually emphasizing more subtle qualities like style or safety). Finally, we use off-the-shelf LLMs as judges for accessibility and reproducibility; exploring specialized evaluators with stronger reasoning or generative reward models may yield further gains.
@@ Acknowledgments
Appendix A Appendix
A.1 Details of RaR-Medicine dataset
The following is illustrates an example from the RaR-Medicine dataset with more details about statistics and categorical distribution of rubrics listed in Table 4 and 6.
RaR-Medicine Example Question. A 50-year-old male patient weighs 65 kg with a pH of 7.05, PCO2 of 15 mmHg, HCO3 of 5 mEq/L, and a base deficit of -40 mEq/L. How much sodium bicarbonate should be administered in the first 4 hours to correct his metabolic acidosis? Reference Answer. To address the severe metabolic acidosis in this 50-year-old patient (weight 65 kg) with a pH of 7.05, it is crucial to approach treatment with caution to avoid complications from overcorrection. The base deficit is -40 mEq/L, which indicates a significant acid-base disturbance. Typically, the formula for bicarbonate administration is: $\text{Bicarbonate required}=\text{Base Deficit}(\text{mEq/L})\times\text{Body Weight}(\text{kg})\times 0.3$ Applying the values: $40\,\text{mEq/L}\times 65\,\text{kg}\times 0.3=780\,\text{mEq}$ However, it is standard practice to administer a partial correction initially, aiming for safety and gradual improvement. Therefore, administering about 150 mEq of sodium bicarbonate over the first 4 hours is a prudent approach. This amount aligns with safe clinical guidelines for initial treatment in severe metabolic acidosis, ensuring the patient starts on the path to stabilization without risking harm from rapid correction. Rubric (checklist and weights). Bicarbonate Calculation (w=5): Essential Criteria: The response must correctly identify and apply the formula (Base Deficit × Body Weight × 0.3) to determine the bicarbonate requirement. Safe Dosing Recommendation (w=5): Essential Criteria: The response must state a clear recommendation of administering about 150 mEq of sodium bicarbonate over the first 4 hours. Partial Correction Justification (w=4): Important Criteria: The response should explain that only a partial correction is administered initially to avoid complications from rapid overcorrection. Step-by-Step Calculation (w=3): Important Criteria: The response must detail the calculation steps, showing that 40 mEq/L × 65 kg × 0.3 equals 780 mEq, before noting the adjusted dose for safety. Base Deficit Interpretation (w=2): Optional Criteria: The response may mention that the base deficit of -40 mEq/L indicates severe metabolic acidosis requiring cautious treatment. Patient Data Accuracy (w=3): Important Criteria: The response must accurately incorporate the patient’s weight of 65 kg and his critical pH, PCO2, and HCO3 values into the explanation. Avoid Overcorrection Risk (w= -1): Pitfall Criteria: Does not mention the risks associated with rapid overcorrection of metabolic acidosis if only the full calculated bicarbonate amount is administered.
@@ A.1 Details of RaR-Medicine dataset
Table 4: Aggregate statistics for the RaR-Medicine dataset (train and validation) dataset.
@@
Important 52,748 34.1
Essential 47,584 30.7
Optional 34,261 22.1
Pitfall 20,215 13.1
@@
Total examples 20,166 100.0
Medical Diagnosis 10,147 50.3
Medical Treatment 3,235 16.0
Medical Knowledge 2,557 12.7
Medical Diag. and Mngmnt 2,033 10.1
@@
A.2 Details of RaR-Science dataset
This section shows an illustrative example from the RaR-Science dataset with more details about statistics and categorical distibution of rubrics listed in Table 7, 8, 9.
RaR-Science Example Question. Determine the solubility of boric acid ($\mathrm{H_{3}BO_{3}}$) in ethanol ($\mathrm{C_{2}H_{5}OH}$) compared to its solubility in benzene ($\mathrm{C_{6}H_{6}}$), considering the principles of ’likes dissolve likes’ and the role of $K_{sp}$ values. Explain your reasoning and provide examples of how immulcifiers could affect the solubility of substances in different solvents. Reference Answer. Boric acid is more soluble in ethanol than in benzene. Rubric (checklist and weights). Correct Solubility Direction (w=5): Essential Criteria: The response must clearly identify that boric acid is more soluble in ethanol than in benzene. Polarity Principle (w=5): Essential Criteria: The answer should explain how the ’like dissolves like’ principle applies by contrasting the polar nature of ethanol with the non-polar character of benzene. Ksp Context (w=4): Important Criteria: The response should account for the role of Ksp values, discussing their typical relevance to solubility even though boric acid is a covalent compound rather than an ionic one. Immulcifier Explanation (w=4): Important Criteria: The answer should explain how immulcifiers could modify solubility, providing an example of their effect on solvation in different solvents. Chemical Properties (w=4): Important Criteria: The response should analyze the inherent chemical properties of boric acid and the solvents to justify the observed solubility differences. Avoid Ionic Assumptions (w= -1): Pitfall Criteria: The answer must not incorrectly assume that Ksp values for ionic compounds directly determine the solubility of a covalent acid such as boric acid. Enhanced Detail (w=2): Optional Criteria: The response may include additional examples or a concise explanation of solvation dynamics to further illustrate how solubility is influenced.